OpenStudy (anonymous):

Solve the system of equations by graphing. -1/3x + y = –1 y = 4 + 1/3x

OpenStudy (tkhunny):

Well, you should graph those two lines and see where they intersect. You may wish to put the first in Slope-Intercept form. It may be easier, that way.

OpenStudy (anonymous):

i have to solve the equations to do that. and i confused myself doing that

OpenStudy (anonymous):

ughh i need help..

OpenStudy (anonymous):

At glance, there is no solution. I can see The 1st equation: 1/3x = y + 1 The 2nd equation: 1/3x = y - 4 --> y + 1 = y - 4 (Impossible) So we don't have any solution, even with graphs If you need to graph these two equations, see the attached file.

OpenStudy (anonymous):

i have to put each equation into y=mx+b form first and im confused on how to do that since i havent done it in a while